OpenStudy (anonymous):

find the exact value of the given trigonometric function cos-11π/4

OpenStudy (anonymous):

cos(-11pi/4)=cos(-11pi/4+2pi)=cos(-3pi/4)=cos(3pi/4)=-cos(pi/4)=-1/sqrt(2) \[-\frac{1}{\sqrt{2}}\]